OpenStudy (chriss):

Side AB of square ABCD is extended to point P so that BP = 2(AB). With M at the midpoint of DC, BM intersects AC at point Q. Also PQ intersects BC at point R. Use Menelaus's theorem to find the numerical value of CR/RB.

OpenStudy (chriss):

I found this solution to this but I don't understand a piece of it. Consider triangle ABC, by Menelaus Theorem, \[\frac{ AQ }{QC } *\frac{ CR }{ RB }*\frac{ PB }{ AP }=1\] \[\frac{ 2 }{ 1 }*\frac{ CR }{ RB }*\frac{ 2 }{ 3 }=1\] \[\frac{ 4 }{ 3 }*\frac{ CR }{ RB }=1\] \[\frac{ CR }{ RB }=\frac{ 3 }{ 4 }\] What I don't understand is where they got 2/1 for AQ/QC. Could someone help me shed some light on that?